Hula Enterprises is considering a new project to produce solar water heaters. The finance manager wishes to find an appropriate risk adjusted discount rate for the project. The (equity) beta of Hot Water, a firm currently producing solar water heaters, is 1.4. Hot Water has a debt to total value ratio of 0.6. The expected return on the market is 0.12, and the riskfree rate is 0.08. Suppose the corporate tax rate is 35 percent. Assume that debt is riskless throughout this problem. (Round your answers to 2 decimal places. (e.g., 0.16))

a. The expected return on the unlevered equity (return on asset, R0) for the solar water heater project is  %.
b. If Hula is an equity financed firm, the weighted average cost of capital for the project is  %.
c. If Hula has a debt to equity ratio of 2, the weighted average cost of capital for the project is  %.
d. The finance manager believes that the solar water heater project can support 45 cents of debt for every dollar of asset value, i.e., the debt capacity is 45 cents for every dollar of asset value. Hence she is not sure that the debt to equity ratio of 2 used in the weighted average cost of capital calculation is valid. Based on her belief, the appropriate debt ratio to use is  %. The weighted average cost of capital that you will arrive at with this capital structure is  %.
Answer #1
a
D/A =0.6
D/E=D/(A-D)=0.6/(1-0.6)=1.5
As per CAPM
expected return = risk-free rate + beta * (expected return on the market - risk-free rate)
Expected return% = 8 + 1.4 * (12 - 8)
Expected return% = 13.6
Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
13.6 = Unlevered cost of equity+1.5*(Unlevered cost of equity-8)*(1-0.35)
Unlevered cost of equity = 10.84
b
For unlevered firm WACC = unlevered cost of equity=10.84%
c
Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
Levered cost of equity = 10.84+2*(10.84-8)*(1-0.35)
Levered cost of equity = 14.53
After tax rate = Cost of debt * (1-Tax rate)
After tax rate = 8 * (1-0.35)
After tax rate = 5.2
D/A = D/(E+D)
D/A = 2/(1+2)
=0.66667
Weight of equity = 1-D/A
Weight of equity = 1-0.66667
W(E)=0.33333
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=5.2*0.66667+14.53*0.33333
WACC=8.31%
d
D/A =0.45
D/E=D/(A-D)=0.45/(1-0.45)=0.8182
Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
Levered cost of equity = 10.84+0.81818*(10.84-8)*(1-0.35)
Levered cost of equity = 12.35
After tax rate = Cost of debt * (1-Tax rate)
After tax rate = 8 * (1-0.35)
After tax rate = 5.2
Weight of equity = 1-0.45
W(E)=0.55
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=5.2*0.45+12.35*0.55
WACC=9.13%